OpenStudy (anonymous):

If f(x)=2x+1 and G(x) x-1/2 then f(g(x))

jimthompson5910 (jim_thompson5910):

is g(x) the function \[\large g(x) = x - \frac{1}{2}\] OR is g(x) the function \[\large g(x) = \frac{x-1}{2}\]

OpenStudy (anonymous):

The second one

jimthompson5910 (jim_thompson5910):

First you start off with f(x) \[\large f(x)=2x+1\] then you replace each copy of x with g(x) like so \[\large f(g(x))=2(g(x))+1\] Afterwards, you replace g(x) on the right side with (x-1)/2 like this \[\large f(g(x))=2\left(\frac{x-1}{2}\right)+1\] The '2's will cancel leaving you with \[\large f(g(x))=x-1+1\] and then you simplify to get this final answer \[\large f(g(x))=x\]

jimthompson5910 (jim_thompson5910):

Hopefully all that makes sense